OpenStudy (anonymous):

What is the solution to the equation? -8x + 9x - 21=-12-16 PLEASE HELP ME!!!

pooja195 (pooja195):

Start by combining like terms

pooja195 (pooja195):

-8x + 9x - 21=-12-16 -8+9=?

OpenStudy (anonymous):

1